Consider the following probability distribution: x: 0 1 2 3 4
p(x): 0.1 0.4 0.3 0.1 0.1
(a) Find the mean of the distribution.
(b) Find the variance.
(c) Find the standard deviation.

Answer by Boreal(15235)

E(X)=mean=0*0.1+1*0.4+2*0.3+3*0.1+4*0.1=1.7
V(X)=1.7^2*0.1+0.7^2*0.4+0.3^2*0.3+1.3^2*0.1+2.3^2*0.1, or different from mean squared * probability and add, which is 0.289+0.196+0.027+0.169+0.529=1.21
Std is sqrt (V)=1.1
mean 1.7, variance 1.21, sd 1.1